Sue travels by car from one city to another. She drives for 23.0 min at 60.0 km/h, 33.0 min at 37.0 km/h, and 30.0 min at 56.0 km/h, and she spends 10.0 min eating lunch.
Determine the average speed for the trip
1
Expert's answer
2012-08-30T09:38:25-0400
The total distance traveled is

L = 60*(23/60) + 37*(33/60) + 56*(30/60) = 71.35 km.

The total time spent is

T = 23 + 33 + 30 + 10 = 96 m = 96/60 h = 1.6 h.

So, the average speed is

V = 71.35 / 1.6 = 44.59375 km/h.